# DPoser: Diffusion Model as Robust 3D Human Pose Prior

### 09 Dec 2023 v1

@@ Abstract
Modeling human pose is a cornerstone in applications from human-robot interaction to augmented reality, yet crafting a robust human pose prior remains a challenge due to biomechanical constraints and diverse human movements. Traditional priors like VAEs and NDFs often fall short in realism and generalization, especially in extreme conditions such as unseen noisy poses. To address these issues, we introduce DPoser, a robust and versatile human pose prior built upon diffusion models. Designed with optimization frameworks, DPoser seamlessly integrates into various pose-centric applications, including human mesh recovery, pose completion, and motion denoising. Specifically, by formulating these tasks as inverse problems, we employ variational diffusion sampling for efficient solving. Furthermore, acknowledging the disparity between the articulated poses we focus on and structured images in previous research, we propose a truncated timestep scheduling to boost performance on downstream tasks. Our exhaustive experiments demonstrate DPoser’s superiority over existing state-of-the-art pose priors across multiple tasks. Corresponding authors: Yulun Zhang and Haoqian Wang.
Figure 1: An overview of DPoser’s versatility and performance across multiple pose-related tasks. Built on diffusion models, DPoser serves as a robust and adaptable pose prior. Shown are applications in (a) pose generation, (b) human mesh recovery, (c) motion denoising, and (d) pose completion. DPoser consistently outstrips existing priors like VPoser [35] in performance benchmarks.
@@ 1 Introduction
The accurate modeling of human pose is an essential building block in a variety of applications, ranging from human-robot interaction to augmented and virtual reality experiences. Many real-world applications rely on a prior distribution of valid human poses to perform tasks like body model fitting, motion capture, and gesture recognition. Existing methods use Gaussian Mixture Models (GMMs) [34], Variational Autoencoders (VAEs) [35], or Neural Distance Fields (NDFs) [36] to model the complex human pose prior. However, these models have limitations. GMMs can generate implausible poses due to their unbounded nature, VAEs are constrained by Gaussian assumptions, and NDFs exhibit limited generalization capabilities, requiring additional training strategies for extreme poses. These shortcomings underscore the need for a more robust and reliable human pose prior, a gap that this work aims to fill.
@@
Diffusion models [10, 8, 12, 14] have recently gained traction for their prowess in capturing complex, high-dimensional data distributions and enabling versatile sampling techniques. Although they have been employed in specialized applications like generating realistic human motion sequences [45, 46] or serving as multi-hypothesis pose estimators based on 2D inputs [41, 43], their utility has generally been constrained. These models are designed for generation or tailored to work with specific types of input data, which limits their applicability in broader contexts. Thus far, the potential of diffusion models as a universal human pose prior remains largely untapped, and effective optimization methods for diverse applications remain unanswered.
@@
In this work, we propose DPoser, a novel approach that leverages time-dependent denoiser learned from expansive motion capture datasets to construct a robust human pose prior. Specifically, by viewing various pose-centric tasks as inverse problems, we propose to employ variational diffusion sampling techniques [18] to incorporate DPoser as a regularization term within the existing optimization framework e.g., SMPLify [34]. Furthermore, our experimental investigations into the generative aspects of articulated human poses reveal that the major information during diffusion is concentrated on the “clean” end of the perturbed data, as opposed to the entire noise-to-clean trajectory. Motivated by this insight, we introduce a truncated annealed timestep scheduling for optimization, which demonstrates superior efficiency over conventional random and full annealed schedules across multiple downstream tasks.
In summary, our main contributions of this work can be summarized as follows:
@@
  • We introduce DPoser, a novel framework based on diffusion models to craft a robust and flexible human pose prior, geared for seamless integration across diverse pose-related applications via test-time optimization.
  • Utilizing a fresh perspective on inverse problems, we employ variational diffusion sampling during test-time optimization and introduce truncated timestep scheduling, showing superior efficiency over traditional methods.
  • Through extensive experiments and ablation studies, we establish that DPoser not only outshines state-of-the-art (SOTA) pose priors in a variety of applications but also provides valuable insights into training techniques such as suitable rotation representations for peak performance.
2 Related works
@@ 2.1 Human Pose Priors
Human body models such as SMPL [33] serve as powerful tools for parameterizing both pose and shape, thereby offering a comprehensive framework for describing a wide array of human gestures. Within the SMPL model, body poses are captured using rotation matrices or joint angles linked to a kinematic skeleton. Adjusting these parameters enables the representation of a diverse range of human actions. However, due to inherent biomechanical constraints, plausible poses inherently reside in a lower-dimensional manifold.
Numerous approaches [34, 35, 36] have been proposed to model this intrinsic pose prior. Classical generative models like GMMs, VAEs [16], and GANs [17] have been employed to capture the complex underlying distribution, demonstrating significant improvements in downstream tasks such as human mesh recovery [49, 50]. Additionally, some research has explored task-specific conditional pose priors, augmenting the model with additional inputs like image features [42, 48], 2D joint coordinates [43], or preceding poses in motion sequences [40, 39]. Our work adopts the unconditional pose prior framework. We train DPoser on large-scale motion capture datasets without auxiliary inputs like paired images or captions, to ensure its applicability across a broad spectrum of pose-related tasks.
@@ 2.2 Diffusion Models for Pose-centric Tasks
Diffusion models [6, 8, 10, 11] have emerged as powerful tools for capturing intricate data distributions, aligning particularly well with the demands of multi-hypothesis estimation in ambiguous human poses. Notable works include DiffPose [41], which leverages a Gaussian Mixture Model-guided forward diffusion process [58] and employs a Graph Convolutional Network (GCN) [57] architecture conditioned on 2D pose sequences for 3D pose estimation by learned reverse process (i.e., generation). In a similar vein, DiffusionPose [42] and GFPose [43] employ the generation-based pipeline but take different approaches in conditioning. ZeDO [44] focuses on 2D-to-3D pose lifting, while works like Diff-HMR [48] and DiffHand [47] extend to estimating SMPL parameters and hand mesh vertices. BUDDI [62] stands out for using diffusion models to capture the joint distribution of interacting individuals and leveraging SDS loss [27, 28] for test-time optimization.
Our work, DPoser, parallels BUDDI in similar optimization implementation. However, we distinguish ourselves by introducing a wider perspective of inverse problems and proposing advanced timestep scheduling based on characteristics of human poses. Moreover, DPoser is designed to work as a flexible regularization term, making it adaptable to diverse pose-centric tasks. It should also be noted that unlike most works [44, 42, 43, 41] that focus on joint positions, DPoser targets the more complex issue of joint rotations, adding an extra layer of challenge due to the intricacy of rotation representations.
3 Preliminary: Score-based Diffusion Models
@@
$\mathrm{d}\mathbf{x}=\mu(t)\mathbf{x}\mathrm{d}t+g(t)\mathrm{d}\mathbf{w}.$ (1)
Here, $\mu(t)\mathbf{x}\in\mathbb{R}^{n}$ and $g(t)\in\mathbb{R}$ represent the drift and diffusion coefficients, while $\mathbf{w}$ is a standard Wiener process.
@@
The so-called score function [4], $\nabla_{\mathbf{x}_{t}}\log p_{t}\left(\mathbf{x}_{t}\right)$, serves as an unknown term in Eq. (2) and can be approximated by a neural network parameterized as $\epsilon_{\theta}(\mathbf{x}_{t};t)\approx-\sigma_{t}\nabla_{\mathbf{x}_{t}}\log p_{t}\left(\mathbf{x}_{t}\right)$ [^1] This parameterization is obtained from the deep connection between the noise prediction in diffusion models and score function estimation in score-based models. We provide a brief recap in the Appendix.. To learn the score functions, employing denoising score matching techniques [2], we first introduce noise to the data points as per:
@@
Subsequently, we train the time-dependent noise predictor $\epsilon_{\theta}(\mathbf{x}_{t};t)$ using an L2-loss defined as [10]:
$\mathbb{E}_{\mathbf{x}_{0}\sim p_{\mathrm{data}},\epsilon\sim\mathcal{N}(\mathbf{0},\mathbf{I}),t\sim\mathcal{U}[0,1]}\left[w(t)||\epsilon-\epsilon_{\theta}(\mathbf{x}_{t};t)||_{2}^{2}\right],$ (4)
where $w(t)$ denotes a positive weighting function. Upon successful training, the score functions can be estimated and used to solve the reverse SDE (Eq. (2)). Through techniques like Euler-Maruyama discretization, we can generate novel samples by simulating the reverse SDE.
4 Methods
4.1 Learning Pose Prior with Diffusion Models
We propose to model pose prior based on the SMPL body model [33], which can be viewed as a differentiable function $[J,V]=M(\theta,\beta)$ that maps body joint angles $\theta\in\mathbb{R}^{3\times 21}$ and shape parameters $\beta\in\mathbb{R}^{10}$ to mesh vertices $V\in\mathbb{R}^{3\times 6890}$ and joint positions $J\in\mathbb{R}^{3\times 22}$. Our target is to model the distribution of joint angles $p(\theta)$. To achieve this, we train DPoser on mocap datasets, employing sub-VP SDEs introduced in ScoreSDE [8], defined as:
$\mathrm{d}\mathbf{x}=-\frac{1}{2}\beta(t)\mathbf{x}\mathrm{d}t+\sqrt{\beta(t)(1-e^{-2\int_{0}^{t}\beta(s)\mathrm{d}s})}\mathrm{d}\mathbf{w},$ (5)
where $\beta(t)$ denotes linear scheduled noise scales. The coefficients needed in Eq. (3) can be obtained as $\alpha_{t}=e^{-\frac{1}{2}\int_{0}^{t}\beta(s)\mathrm{d}s},\sigma_{t}=[1-e^{-\int_{0}^{t}\beta(s)\mathrm{d}s}]^{2}$. With the constructed forward process (Eq. (3)) to sample $\mathbf{x}_{t}$, we then apply the objective in Eq. (4) to train the noise predictor $\epsilon_{\theta}(\mathbf{x}_{t};t)$ with weights $w(t)=\sigma_{t}^{2}$ as suggested in ScoreSDE [8].
In our experiments, we find the selection of rotation representations, such as axis-angle and 6D rotations [59], to be crucial for optimal performance when coupled with suitable normalization. Despite experimenting with incorporating body models via auxiliary losses, as in prior work [35, 39], we note that it imposes a computational burden without evident gains in downstream task performance. Further methodological nuances are detailed in the Appendix.
@@ 4.2 Optimization Leveraging Diffusion Priors
Figure 2: Overview of the DPoser Methodology. Panel (a) presents three tasks: human mesh recovery, pose completion, and motion denoising, with omissions like camera optimization and bending loss for clarity. Panel (b) demonstrates the DPoser regularization process, introducing noise to the current pose and applying a one-step denoiser to achieve a denoised pose. $L_{\text{DPoser}}$ is computed between the denoised and current pose. Panel (c) outlines the optimization process from initial to current and fitted poses via loss minimization.
The acquired score functions or noise predictors, denoted as $\epsilon_{\theta}(\mathbf{x}_{t};t)$, permit the direct generation of plausible poses through Eq. (2). Yet, the broader integration of diffusion priors into general optimization frameworks remains an open avenue. We address this by reframing pose-related tasks as inverse problems and applying variational diffusion sampling techniques [18] for efficient resolution.
@@
$\mathbf{y}=\mathcal{A}(\mathbf{x}_{0})+\mathbf{n},\quad\mathbf{y},\mathbf{n}\in\mathbb{R}^{d},~{}\mathbf{x}_{0}\in\mathbb{R}^{n},$ (6)
where $\mathcal{A}$ symbolizes the measurement operator and $\mathbf{n}$ constitutes noise, assumed to be white Gaussian $\mathcal{N}(\mathbf{0},\sigma_{n}^{2}\mathbf{I})$ in our context. As an illustration, inverse kinematics can be considered as inferring joint rotations (or body poses in SMPL [33]) from observed joint positions. Moreover, the 2D-3D lifting task can be modeled as a linear inverse problem, with $\mathcal{A}$ representing a camera projection model. To recover the original signal $\mathbf{x}_{0}$, in the Bayesian framework, our target is to sample from the posterior distribution $p\left(\mathbf{x}_{0}\mid\mathbf{y}\right)$.
Diffusion-based inverse problem solving. Various techniques [26, 21, 19, 20, 22, 18] have been explored to simulate this posterior sampling process based on diffusion priors. For some noiseless linear inverse problems, such as completion, alternating projections onto measurement subspaces are commonly employed [8, 21, 20]. For instance, in hand pose completion, a standard iterative generation process using diffusion models can be applied, replacing known parts of the pose at each iteration.
However, such projections are inconvenient to build general optimization frameworks concerned in our work. To navigate these challenges, we adopt variational diffusion sampling [18], efficient in computation and straightforward in implementation. Specifically, it employs a variational distribution $q\left(\mathbf{x}_{0}\mid\mathbf{y}\right):=\mathcal{N}(\mu,\sigma^{2}\mathbf{I})$ and aims to minimize the Kullback-Leibler (KL) divergence between this variational distribution and the true posterior, mathematically expressed as $KL\big{(}q\left(\mathbf{x}_{0}\mid\mathbf{y}\right)\parallel p\left(\mathbf{x}_{0}\mid\mathbf{y}\right)\big{)}$.
Informed by existing theorems concerning the likelihood properties of diffusion models [9], and under the assumption of zero variance ($\sigma\approx 0$), the optimization problem of seeking $\mathbf{x}_{0}$ (i.e., $\mu$) can be formulated as minimizing [18]:
$\|\mathbf{y}-\mathcal{A}(\mathbf{x}_{0})\|^{2}+w_{t}(\mathtt{sg}[\epsilon_{\theta}(\mathbf{x}_{t};t)-\epsilon])^{\top}\mathbf{x}_{0},$ (7)
@@
where $w_{t}$ denotes the loss weights and $\epsilon$ is sampled from the standard Gaussian distribution. Here, $\mathtt{sg}$ signifies the stopped-gradient operator, indicating that backpropagation through the trained diffusion models is not required. The optimization procedure initiates by selecting a timestep $t$ and applying a perturbation to the target $\mathbf{x}_{0}$ as per Eq. (3), resulting in $\mathbf{x}_{t}$. Subsequently, the gradients $[\epsilon_{\theta}(\mathbf{x}_{t};t)-\epsilon]$ are applied to the optimization variable $\mathbf{x}_{0}$. In a nutshell, this framework [18] provides a flexible yet robust strategy for employing diffusion priors in generic optimization problems, serving as a cornerstone for our work.
@@
DPoser regularization. To shed more light on the working mechanism, we reformulate the regularization term as [^2] We detail another perspective of Score Distillation Sampling (SDS) [27, 28] to understand our DPoser regularization in the Appendix.:
@@
$\mathbf{\hat{x}}_{0}(t)$ $=\frac{\mathbf{x}_{t}-\sigma_{t}\epsilon_{\theta}(\mathbf{x}_{t};t)}{\alpha_{t}}.$ (9)
Here, $\mathbf{\hat{x}}_{0}(t)$ serves as a one-step denoised estimate, leveraging the trained diffusion model $\epsilon_{\theta}(\mathbf{x}_{t};t)$. We show that the gradient of our DPoser regularization aligns with Eq. (7).
@@
$=2w_{t}(\frac{\mathbf{x}_{t}-\sigma_{t}\epsilon}{\alpha_{t}}-\frac{\mathbf{x}_{t}-\sigma_{t}\epsilon_{\theta}(\mathbf{x}_{t};t)}{\alpha_{t}})$
$=2w_{t}\frac{\sigma_{t}}{\alpha_{t}}(\epsilon_{\theta}(\mathbf{x}_{t};t)-\epsilon)$
$\propto(\epsilon_{\theta}(\mathbf{x}_{t};t)-\epsilon).$ (10)
Therefore, our DPoser regularization $L_{\mathrm{DPoser}}$ can be implemented as an L2-loss. In this view, during optimization, DPoser works by pulling the current pose $\mathbf{x}_{0}$ closer to the learned plausible pose distribution through denoising.
DPoser across pose-related tasks. Capitalizing on DPoser’s inherent flexibility, we adapt it to various human pose tasks, with a spotlight on human mesh recovery as illustrated in Fig. 2. For an exhaustive examination of DPoser’s utility across different pose-related applications including pose completion and motion denoising, we direct the reader to our Appendix.
By adapting the objective function from the SMPLify framework [34], we supplant the originally used GMM prior with our own DPoser regularization term, $L_{\mathrm{DPoser}}$, while also forgoing the intricate interpenetration error term. The optimization objective, incorporating both pose $\theta$ and shape $\beta$ parameters in SMPL [33], is articulated as:
@@
Herein, the reprojection loss $L_{J}$ serves as the data fidelity term, mathematically expressed as:

### 23 Mar 2024 v2

@@ Abstract
This work targets to construct a robust human pose prior. However, it remains a persistent challenge due to biomechanical constraints and diverse human movements. Traditional priors like VAEs and NDFs often exhibit shortcomings in realism and generalization, notably with unseen noisy poses. To address these issues, we introduce DPoser, a robust and versatile human pose prior built upon diffusion models. DPoser regards various pose-centric tasks as inverse problems and employs variational diffusion sampling for efficient solving. Accordingly, designed with optimization frameworks, DPoser seamlessly benefits human mesh recovery, pose generation, pose completion, and motion denoising tasks. Furthermore, due to the disparity between the articulated poses and structured images, we propose truncated timestep scheduling to enhance the effectiveness of DPoser. Our approach demonstrates considerable enhancements over common uniform scheduling used in image domains, boasting improvements of 5.4%, 17.2%, and 3.8% across human mesh recovery, pose completion, and motion denoising, respectively. Comprehensive experiments demonstrate the superiority of DPoser over existing state-of-the-art pose priors across multiple tasks. Corresponding authors: Yulun Zhang and Haoqian Wang.
Keywords:
Human Pose Prior, Diffusion Model
@@ 1 Introduction
Accurate modeling of human pose is a fundamental research topic that can benefit various applications, from human-robot interaction to augmented and virtual reality experiences. Many real-world applications rely on a prior distribution of valid human poses to perform tasks like body model fitting, motion capture, and gesture recognition. The complexity of human biomechanics, coupled with the extensive kinematic variability in movement patterns, presents a significant challenge in constructing a robust and realistic human pose prior.
Previous efforts to model human pose prior have mainly employed techniques such as Gaussian Mixture Models (GMMs) [1], Variational Autoencoders (VAEs) [37], and Neural Distance Fields (NDFs) [49]. Each technique, however, faces its own set of limitations. GMMs, for instance, might lead to the generation of implausible poses due to their unbounded nature. VAEs, restricted by their Gaussian assumptions, tend to generate average poses that may not accurately capture the full spectrum of human actions. Meanwhile, NDFs have shown promise in 3D surface modeling but struggle with generalizing across the complex, high-dimensional landscape of human pose manifolds. These limitations highlight a pressing need for a more comprehensive and dependable approach to modeling human pose priors, an endeavor this work seeks to address.
@@
Recently, Diffusion models [16, 48, 11, 20] have gained traction for their prowess in capturing complex, high-dimensional data distributions and enabling versatile sampling techniques. Their application has been seen in generating lifelike human motion sequences [56, 42] and functioning as multi-hypothesis pose estimators from 2D inputs [17, 8]. However, these models are designed for specific generation tasks or tailored to work with conditional input data, which limits their applicability in broader contexts. The potential of diffusion models as a universal human pose prior remains largely untapped, and effective optimization methods for diverse tasks remain unanswered.
Figure 1: An overview of DPoser’s versatility and performance across multiple pose-related tasks. Built on diffusion models, DPoser serves as a robust and adaptable pose prior. Shown are scenarios in (a) pose generation, (b) human mesh recovery, (c) motion denoising, and (d) pose completion. DPoser consistently outstrips existing priors like VPoser [37] in performance benchmarks.
@@
In this work, we propose DPoser, a novel approach that leverages time-dependent denoiser learned from expansive motion capture datasets to construct a robust human pose prior. We regard various pose-centric tasks as inverse problems and suggest the integration of DPoser via variational diffusion sampling techniques [33] as a regularization component within optimization frameworks like SMPLify [1]. Furthermore, our investigations reveal that significant pose-related information during diffusion is predominantly located at the latter stages of the diffusion trajectory. This revelation inspired us to develop a novel truncated timestep scheduling strategy for optimization. Our method outperforms the standard uniform scheduling, showing gains of 5.4%, 17.2%, and 3.8% in human mesh recovery, pose completion, and motion denoising, respectively.
In summary, our main contributions are as follows:
@@
  • We introduce DPoser, a novel framework based on diffusion models to craft a robust and flexible human pose prior, geared for seamless integration across diverse pose-related tasks via test-time optimization.
  • We analyze the impact of diffusion timesteps in the pose domain and propose truncated scheduling for more efficient optimization.
  • Through extensive experiments, we establish that DPoser outshines state-of-the-art (SOTA) pose priors in a variety of downstream tasks.
2 Related Work
@@ 2.1 Human Pose Priors
Human body models such as SMPL [31] serve as powerful tools for parameterizing both pose and shape, thereby offering a comprehensive framework for describing human gestures. Within the SMPL model, body poses are captured using rotation matrices or joint angles linked to a kinematic skeleton. Adjusting these parameters enables the representation of a diverse range of human actions. Nonetheless, feeding unrealistic poses into these models can result in non-viable human figures, primarily because plausible human poses are confined within a complex, high-dimensional manifold due to biomechanical constraints.
Various strategies [1, 37, 49, 9] have been put forward to build human pose priors. Generative frameworks like GMMs, VAEs [22], and Generative Adversarial Networks (GANs) [13] have shown promise in encapsulating the multifaceted pose distribution, facilitating advancements in tasks like human mesh recovery [19, 12]. Further, some studies have delved into conditional pose priors tailored to specific tasks, incorporating extra information such as image features [39, 3], 2D joint coordinates [8], or sequences of preceding poses [28, 40]. Our initiative leans towards an unconditional pose prior approach, training DPoser on extensive motion capture data without relying on additional inputs like images or text, aiming for a versatile application across various pose-related scenarios.
@@ 2.2 Diffusion Models for Pose-centric Tasks
Diffusion models [47, 48, 16, 44] have emerged as powerful tools for capturing intricate data distributions, aligning particularly well with the demands of multi-hypothesis estimation in ambiguous human poses. Notable works include DiffPose [17], which leverages a Gaussian Mixture Model-guided forward diffusion process [36] and employs a Graph Convolutional Network (GCN) [23] architecture conditioned on 2D pose sequences for 3D pose estimation by learned reverse process (i.e., generation). In a similar vein, DiffusionPose [39] and GFPose [8] employ the generation-based pipeline but take different approaches in conditioning. Further, ZeDO [18] concentrates on 2D-to-3D pose lifting, while Diff-HMR [3] and DiffHand [24] explore estimating SMPL parameters and hand mesh vertices, respectively. BUDDI [34] stands out for using diffusion models to capture the joint distribution of interacting individuals and leveraging SDS loss [38, 52] for optimization during testing phases.
While DPoser shares a similar optimization implementation with BUDDI, it sets itself apart by introducing a wider perspective of inverse problems and equipping an innovative timestep scheduling strategy tailored to the characteristics of human poses. Unlike other approaches [18, 39, 8, 17] that primarily focus on 3D location-based representation, DPoser takes on the more demanding task of modeling SMPL-based rotation pose representation. This adds complexity due to the intricacies involved in representing rotations, positioning DPoser as a more versatile solution within the realm of pose-centric tasks.
3 Methods
3.1 Preliminary: Score-based Diffusion Models
@@
$\mathrm{d}\mathbf{x}=\mu(t)\mathbf{x}\mathrm{d}t+g(t)\mathrm{d}\mathbf{w},$ (1)
where $\mu(t)\mathbf{x}\in\mathbb{R}^{n}$ and $g(t)\in\mathbb{R}$ represent the drift and diffusion coefficients, while $\mathbf{w}$ is a standard Wiener process.
@@
The so-called score function [29], $\nabla_{\mathbf{x}_{t}}\log p_{t}\left(\mathbf{x}_{t}\right)$, serves as an unknown term in Eq. (2) and can be approximated by a neural network parameterized as $\epsilon_{\phi}(\mathbf{x}_{t};t)\approx-\sigma_{t}\nabla_{\mathbf{x}_{t}}\log p_{t}\left(\mathbf{x}_{t}\right)$ [^1] This parameterization is obtained from the deep connection between the noise prediction in diffusion models and score function estimation in score-based models. We provide a brief recap in the Appendix.. To learn the score functions, employing denoising score matching techniques [50], we perturb the data points with noise as per:
@@
Subsequently, feeding $\mathbf{x}_{t}$ and $t$ as input, we train the time-dependent noise predictor $\epsilon_{\phi}(\mathbf{x}_{t};t)$ using an L2-loss defined as [16]:
$\mathbb{E}_{\mathbf{x}_{0}\sim p_{\mathrm{data}},\epsilon\sim\mathcal{N}(\mathbf{0},\mathbf{I}),t\sim\mathcal{U}[0,1]}\left[w(t)||\epsilon-\epsilon_{\phi}(\mathbf{x}_{t};t)||_{2}^{2}\right],$ (4)
where $w(t)$ denotes a positive weighting function.
Upon successful training, the score functions can be estimated and used to solve the reverse SDE (Eq. (2)). Through techniques like Euler-Maruyama discretization, we can generate novel samples by simulating the reverse SDE.
3.2 Learning Pose Prior with Unconditional Diffusion Models
SMPL-based pose representation. To build a flexible 3D human pose prior, we propose to utilize the SMPL body model [31], which can be viewed as a differentiable function $[J,V]=M(\theta,\beta)$ that maps body joint angles $\theta\in\mathbb{R}^{3\times 21}$ and shape parameters $\beta\in\mathbb{R}^{10}$ to mesh vertices $V\in\mathbb{R}^{3\times 6890}$ and joint positions $J\in\mathbb{R}^{3\times 22}$. Our target is to model the distribution of joint angles $p(\theta)$.
Training of unconditional diffusion models. To this end, we adopt an unconditional diffusion model to learn the pose representation $\theta$. This approach aligns with a task-agnostic strategy, focusing solely on the distribution of 3D poses. We employ sub-VP SDEs as outlined in [48], which have demonstrated efficacy in sampling quality, for constructing our diffusion model. Specifically, our chosen forward SDE (Eq. (1)) is given by:
$\mathrm{d}\mathbf{x}=-\frac{1}{2}\xi(t)\mathbf{x}\mathrm{d}t+\sqrt{\xi(t)(1-e^{-2\int_{0}^{t}\xi(s)\mathrm{d}s})}\mathrm{d}\mathbf{w},$ (5)
where $\xi(t)$ denotes linear scheduled noise scales. The coefficients needed in Eq. (3) can be obtained as $\alpha_{t}=e^{-\frac{1}{2}\int_{0}^{t}\xi(s)\mathrm{d}s},\sigma_{t}=1-e^{-\int_{0}^{t}\xi(s)\mathrm{d}s}$.
During training, we initiate with a clean data point $\mathbf{x}_{0}$—essentially, our pose representation $\theta$—and introduce noise to generate samples $\mathbf{x}_{t}$ according to the forward process detailed in Eq. (3). Then we apply the objective in Eq. (4) to train the noise predictor $\epsilon_{\phi}(\mathbf{x}_{t};t)$ with weights $w(t)=\sigma_{t}^{2}$ as suggested in [48].
@@ 3.3 Optimization Leveraging Diffusion Priors
The acquired score functions or noise predictors, denoted as $\epsilon_{\phi}(\mathbf{x}_{t};t)$, permit the direct generation of plausible poses through Eq. (2). Yet, the broader integration of diffusion priors into general optimization frameworks remains an open avenue. We address this by reframing pose-related tasks as inverse problems and applying variational diffusion sampling techniques [33] for efficient resolution.
@@
$\mathbf{y}=\mathcal{A}(\mathbf{x}_{0})+\mathbf{n},\quad\mathbf{y},\mathbf{n}\in\mathbb{R}^{d},~{}\mathbf{x}_{0}\in\mathbb{R}^{n},\vspace{-2mm}$ (6)
where $\mathcal{A}$ symbolizes the measurement operator and $\mathbf{n}$ constitutes noise, assumed to be white Gaussian $\mathcal{N}(\mathbf{0},\sigma_{n}^{2}\mathbf{I})$. In the context targeted in this study, $\mathbf{x}_{0}$ always refers to body poses in SMPL [31]. This formulation allows us to approach various pose-centric tasks by adapting $\mathcal{A}$ and interpreting $\mathbf{y}$ accordingly:
  • Pose completion: Here, $\mathcal{A}$ serves as a mask matrix to simulate partially observed poses, with $\mathbf{y}$ being the incomplete pose data.
  • Motion denoising: In this scenario, $\mathcal{A}$ applies SMPL’s forward kinematics, treating $\mathbf{y}$ as the observed noisy 3D joints.
  • Human mesh recovery: $\mathcal{A}$ integrates SMPL’s forward kinematics and camera projection to relate $\mathbf{y}$ to 2D joint observations in images.
The aim is to recover the original signal $\mathbf{x}_{0}$, where, within the Bayesian framework, our objective shifts to sampling from the posterior distribution $p\left(\mathbf{x}_{0}\mid\mathbf{y}\right)$.
Solving inverse problems with diffusion models. Various techniques [14, 21, 6, 7, 45, 33] have been explored to simulate this posterior sampling process based on unconditional diffusion priors $p\left(\mathbf{x}_{0};\phi\right)$. Among them, the sampling-based scheme is widely explored and applied in tasks like image restoration. These methods incorporate the observation information $\mathbf{y}$ into the generation process of $\mathbf{x}_{0}$ through techniques like gradient guidance [6, 7] and back projection [48, 21, 7]. However, such methods rooted in generation are inconvenient for handling diverse pose-related tasks. To navigate these challenges, we adopt variational diffusion sampling [33] to build general optimization frameworks. Specifically, it employs a variational distribution $q\left(\mathbf{x}_{0}\mid\mathbf{y}\right):=\mathcal{N}(\mu,\sigma^{2}\mathbf{I})$ and aims to minimize the Kullback-Leibler (KL) divergence between this variational distribution and the true posterior, mathematically expressed as $KL\big{(}q\left(\mathbf{x}_{0}\mid\mathbf{y}\right)\parallel p\left(\mathbf{x}_{0}\mid\mathbf{y}\right)\big{)}$. Further, under the assumption of zero variance ($\sigma\approx 0$), the optimization problem of seeking $\mathbf{x}_{0}$ (i.e., $\mu$) can be formulated as minimizing [46, 33]:
$\|\mathbf{y}-\mathcal{A}(\mathbf{x}_{0})\|^{2}+w_{t}(\mathtt{sg}[\epsilon_{\phi}(\mathbf{x}_{t};t)-\epsilon])^{\top}\mathbf{x}_{0},\vspace{-1.5mm}$ (7)
@@
where $w_{t}$ denotes the loss weights and $\epsilon$ is sampled from the standard Gaussian distribution. Here, $\mathtt{sg}$ signifies the stopped-gradient operator, indicating that backpropagation through the trained diffusion models is not required. The optimization procedure initiates by selecting a timestep $t$ and applying a perturbation to the target $\mathbf{x}_{0}$ as per Eq. (3), resulting in $\mathbf{x}_{t}$. Subsequently, the gradients $[\epsilon_{\phi}(\mathbf{x}_{t};t)-\epsilon]$ are applied to the optimization variable $\mathbf{x}_{0}$. In a nutshell, this framework [33] provides a flexible yet robust strategy for employing diffusion priors in generic optimization problems, serving as a cornerstone for our work.
Figure 2: Overview of the DPoser Methodology. Panel (a) presents three tasks: human mesh recovery, pose completion, and motion denoising, with omissions like camera optimization for clarity. Panel (b) demonstrates the DPoser regularization process, introducing noise to the current pose and applying a one-step denoiser to achieve a denoised pose. $L_{\text{DPoser}}$ is computed between the denoised and current pose. Panel (c) outlines the optimization process from initial to fitted poses via loss minimization.
@@
Introducing DPoser regularization. To shed more light on the working mechanism, we propose to reformulate the regularization term as:
@@
$\mathbf{\hat{x}}_{0}(t)$ $=\frac{\mathbf{x}_{t}-\sigma_{t}\epsilon_{\phi}(\mathbf{x}_{t};t)}{\alpha_{t}}.$ (9)
Here, $\mathbf{\hat{x}}_{0}(t)$ functions as a precise one-step denoising prediction using the diffusion model $\epsilon_{\phi}(\mathbf{x}_{t};t)$. This approach effectively encourages the current pose $\mathbf{x}_{0}$ towards a denoised, plausible pose distribution, employing a straightforward L2-loss within the DPoser regularization framework. Further, the theoretical foundation of our regularization demonstrates its alignment with the gradient direction of variational diffusion sampling (Eq. (7)).
@@
$=2w_{t}(\frac{\mathbf{x}_{t}-\sigma_{t}\epsilon}{\alpha_{t}}-\frac{\mathbf{x}_{t}-\sigma_{t}\epsilon_{\phi}(\mathbf{x}_{t};t)}{\alpha_{t}})$
$=2w_{t}\frac{\sigma_{t}}{\alpha_{t}}(\epsilon_{\phi}(\mathbf{x}_{t};t)-\epsilon)$
$\propto(\epsilon_{\phi}(\mathbf{x}_{t};t)-\epsilon).$ (10)
Thus, $L_{\mathrm{DPoser}}$ represents a more intuitive approach to variational diffusion sampling. By incorporating alongside task-specific loss functions, this regularization term enhances the plausibility of the resultant poses.
DPoser across pose-related tasks. DPoser excels in versatility, enabling its seamless application in a spectrum of human pose-related tasks. Its adaptability is especially evident in our human mesh recovery approach, as depicted in Fig. 2. For an exhaustive examination of DPoser’s utility across tasks like pose completion and motion denoising, we direct the reader to our Appendix.
Human mesh recovery aims to deduce the human pose and shape from single-image inputs. In this context, we refine the optimization function derived from the SMPLify framework [1], integrating DPoser as a regularization term, $L_{\mathrm{DPoser}}$, and streamlining the process by omitting the intricate interpenetration error component. The modified optimization objective, engaging both pose $\theta$ and shape $\beta$ parameters from the SMPL model [31], is defined as:
@@
The reprojection loss $L_{J}$, acting as the data fidelity measure, is defined by: